### v3

# XLD: A Cross-Lane Dataset for Benchmarking Novel Driving View Synthesis

Abstract
Comprehensive testing of autonomous systems through simulation is essential to ensure the safety of autonomous driving vehicles. This requires the generation of safety-critical scenarios that extend beyond the limitations of real-world data collection, as many of these scenarios are rare or rarely encountered on public roads. However, evaluating most existing novel view synthesis (NVS) methods relies on sporadic sampling of image frames from the training data, comparing the rendered images with ground-truth images. Unfortunately, this evaluation protocol falls short of meeting the actual requirements in closed-loop simulations. Specifically, the true application demands the capability to render novel views that extend beyond the original trajectory (such as cross-lane views), which are challenging to capture in the real world. To address this, this paper presents a synthetic dataset for novel driving view synthesis evaluation, which is specifically designed for autonomous driving simulations. This unique dataset includes testing images captured by deviating from the training trajectory by $1-4$ meters. It comprises six sequences that cover various times and weather conditions. Each sequence contains $450$ training images, $120$ testing images, and their corresponding camera poses and intrinsic parameters. Leveraging this novel dataset, we establish the first realistic benchmark for evaluating existing NVS approaches under front-only and multicamera settings. The experimental findings underscore the significant gap in current approaches, revealing their inadequate ability to fulfill the demanding prerequisites of cross-lane or closed-loop simulation. Our dataset and code are released publicly on the project page: https://3d-aigc.github.io/XLD.
1 Introduction
Autonomous driving (AD) simulation, which bridges the gap between the real and virtual worlds, is essential for testing and developing autonomous driving software in vehicles [7]. Research indicates that employing effective simulation methods can significantly expedite the evaluation of safety tests for autonomous driving, achieving a speedup of approximately $10^{3}$ to $10^{5}$ times faster than real-world testing [21]. This compelling evidence underscores the importance of leveraging simulation to enhance the efficiency and effectiveness of autonomous driving development. However, the self-driving industry primarily conducts system testing using two approaches: log replay, which involves testing on pre-recorded real-world sensor data, and real-world driving, where new miles are driven to gather additional data for testing purposes [61]. In closed-loop simulation, the vehicle must be free to respond to control commands within the simulation environment rather than strictly following the original trajectory from logs. To promote the rapid advancement of end-to-end autonomous driving systems [64, 24], designing a neural simulator for AD simulation [61, 57, 55], which can render photo-realistic images on novel views for closed-loop simulation and algorithm training, is in high demands. The main scientific problems boil down to the 3D reconstruction [35], and novel view synthesis (NVS) [59, 57], which are also long-standing problems in computer vision and computer graphics. Traditional methods such as [43, 44] have dominated the major deployment of 3D scene reconstruction for a long time. However, those reconstructed scenes cannot be directly used to produce photo-realistic novel views, thus imposing large restrictions on sensor simulations. As a result, the industrial bridges the sim-to-real gap by parametric and procedural modeling technique [50] or human-involved creations. With the recent rapid development of 3D implicit fields, such as neural radiance field (NeRF [37]) and explicit primitive representations, i.e., 3D Gaussian Splatting (3DGS [26]), reconstructing a scene from a collection of images serves as the foundation of end-to-end autonomous driving simulation [23]. These techniques enable the rendering of high-quality and photorealistic images on novel views.
Presently, most approaches evaluate the performance of NVS results by splitting the dataset into training and testing sets. However, this strategy of splitting and sampling leads to an interpolation benchmark, which we argue is insufficient for evaluating whether the trained models can effectively render simulation-ready (i.e., cross-lane) and high-fidelity data for closed-loop simulations. On the contrary, our proposed XLD dataset is a brand-new benchmark that evaluates the synthesis quality on a cross-lane view with additionally captured GT images. Our dataset and benchmark focus on assessing the NVS capability specifically for cameras in cross-lane scenes. The primary objective is to evaluate the performance of cameras in generating accurate and realistic novel views in scenarios involving multiple lanes. Specifically, we introduce the XLD dataset, which encompasses the generation of $150\times 3$ rendered images for each scene. Furthermore, we evaluate novel cross-lane view synthesis by rendering $30$ images with deviations of 00m, $1$m, $2$m, and $4$m from the training trajectory. Using the XLD dataset, we conduct a benchmark of leading methods, which are based on either NeRF or 3DGS, using well-established NVS metrics. Our benchmark results demonstrate that the proposed dataset offers a comprehensive evaluation benchmark tailored specifically to the requirements of closed-loop simulation. Furthermore, our benchmarking results reveal intriguing findings, emphasizing the value of the proposed dataset.
2 Related Work
2.1 Autonomous Driving Simulation
In the past few years, there has been a surge in the use of autonomous driving simulations [31]. These simulators are instrumental in validating planning and control mechanisms, producing educational and evaluative datasets, and significantly cutting down the time needed to perform these functions. The current landscape is dominated by two predominant categories of simulation tools: model-based and data-driven. Model-based simulation platforms, such as PyBullet [18] MuJoCo [49], AirSim [45] and CARLA [19], utilize advanced computer graphics to replicate vehicles and their surroundings. However, the manual effort required to construct these models and program the vehicles’ dynamics can be quite demanding and lengthy. Moreover, the visual output may sometimes fall short of the necessary realism, which can adversely affect the efficacy of perception systems when they are put into operation.
Previously, NVS heavily relied on conventional image processing techniques. For instance, Chaurasia et al.[12] propose using depth synthesis from over-segmented graph structures. At the same time, AADS[30] employs filtered and completed dense depth maps for warping novel view images through image stitching. A data-driven simulation platform VISTA [3, 2] leverages datasets from the real world to create comprehensively labeled and photorealistic simulations. Recently, a wave of innovations has employed the NeRF method to simulate driving perspectives superficially. These new approaches excel in creating photorealistic images and have been shown to surpass traditional view synthesis algorithms in the realm of autonomous driving simulation. Recent advances in neural novel view synthesis significantly accelerate the rapid development of the next-gen driving simulation, which exhibits superior expressiveness and flexibility compared to traditional methods. Our dataset and benchmark are specifically designed for those methods.
2.2 NeRF-based NVS for Driving Simulation
The introduction of neural radiance field (NeRF) revolutionized NVS by incorporating coordinate-based representation within multilayer perceptron (MLP) architectures, leading to significant performance improvements. Building upon NeRF, numerous subsequent works have further adapted these algorithms to fulfill requirements such as efficient training, anti-aliasing rendering, large-scale reconstruction, etc.. InstantNGP [38] proposes using a multi-resolution hash grid with a shallow MLP network to eliminate large MLP networks. Mip-NeRF [4, 5] uses anti-aliased conical frustums instead of rays to reduce objectionable aliasing artifacts, which enables NeRF to represent fine details. Zip-NeRF [6] borrows the ideas from rendering and signal processing that combine Mip-NeRF with InstantNGP. Nerfacto [48] integrates many advantages of existing methods to provide an all-in-one solution for NeRF training. Block-NeRF [47] tackles the reconstruction of large-scale urban scenes by division. To handle dynamics, NSG [41] decomposes dynamic scenes into scene graphs and learns a structured representation. SUDS [51] factorizes a large scene into three hash table data structures, encoding static, dynamic, and far-field radiance fields. MARS [57] is an instance-aware and modular simulator based on NeRF, which models dynamic foreground instances and static background environments separately. UniSim [61] transforms a recorded log into a realistic closed-loop multi-sensor simulation, which incorporates dynamic object priors and utilizes a convolutional network to handle and complete unseen regions. EmerNeRF [60] employs a self-bootstrapping approach to simultaneously capture scene geometry, appearance, motion, and semantics, which enables comprehensive and synchronized modeling of these elements by stratifying scenes into static and dynamic fields. UC-NeRF [15] addresses the challenge of under-calibrated multi-view novel view synthesis through layer-based color correction and virtual warping techniques.
Figure 1: Our datasets encompass six distinct scenes, each involving the vehicle following an on-road trajectory. To generate training data for the cameras and LiDAR sensor, we sample $150$ waypoints along each trajectory. The trajectory is visually emphasized using the color red.
2.3 3DGS-based NVS for Driving Simulation
Motivated by the NeRF-based methods and point-based differentiable rendering [62, 1, 42, 32], 3D Gaussian Splatting (3DGS) [26] opens a new era with the leading advantages in explicit representation and real-time rendering capability. Within a concise timeframe, numerous methods [13, 59] have emerged that focus on road scene reconstruction and NVS by leveraging the 3DGS representation. For instance, PVG [13, 29] introduces periodic vibration-based temporal dynamics to reconstruct dynamic urban scenes. StreetGaussian [59] models the dynamic urban street environment as a collection of point clouds with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. DrivingGaussian [65] uses incremental static 3D Gaussians to represent the scene’s static background. It also employs a composite dynamic Gaussian graph to handle multiple moving objects with LiDAR data. HO-Gaussian [33] introduces a hybrid method that combines radiance fields with 3DGS representation, eliminating the requirement for point initialization in urban scene NVS. HGS-Mapping [56] proposes a hybrid Gaussian representation specifically designed for performing online dense mapping in unbounded large-scale scenes. GaussianPro [16] utilizes priors from reconstructed scene geometries and patch-matching techniques to generate precise Gaussians, leveraging the scene’s existing structure. DC-Gaussian [53] introduces adaptive image decomposition for modeling reflections and occlusions. It incorporates illumination-aware obstruction modeling to handle reflections and occlusions under varying lighting conditions in urban scene novel view synthesis. Our dataset and benchmark specifically focus on evaluating the performance of neural-based driving simulation in NVS, particularly in cross-lane scenarios. A portion of the mentioned methods serve as our baselines, taking into account the code availability. Moreover, generalizable 3D-GS methods such as PixelSplat [11], Mvsplat [14], and GGRt [28] also attempts to synthesize novel images within a well-trained generalizable feed-forward Gaussian networks.
Figure 2: The composition of our training set and testing set. The training set consists of three RGB cameras (’front’, ’left-front’, and ’right-front’) mounted on our vehicle lane. We sample image sequences along trajectories that run parallel to the vehicle’s route for the test set. This encompasses four distinct test trajectories, each offset from the vehicle’s trajectory by 0 meters, 1 meter, 2 meters, and 4 meters, respectively. The sampling interval for the test set is five times the sampling interval used for the training set.
2.4 Datasets in Autonomous Driving
In autonomous driving training and benchmarking, there are many datasets available. For example, KITTI [22]. KITTI-360 [34], vKITTI [8], CityScapes [17], Mapillary [39], ApolloScape [25], Waymo Open Dataset [46]. nuScenes [9], Argoverse [10] and Argoverse 2 [54], BDD100K [63], OpenLane-V2 [52], LiDAR-CS [20], etc.. Those previous works have laid the groundwork for research and development in autonomous driving algorithms. A comprehensive survey of datasets related to autonomous driving refers to [36]. More recently, a newly proposed Hierarchical 3D-GS [27] showcases the scalability of 3D-GS followed by a short scene with parallel trajectories to train and render. However, none of these existing researches have specifically addressed the evaluation of novel view synthesis techniques tailored for autonomous driving simulation, particularly in terms of their ability to meet the high demands of cross-lane capability.
3 Dataset
Distinguishing our dataset from previous datasets (both real-world and synthetic) that typically capture only a single road trajectory, ours additionally captures multiple parallel trajectories. To this end, we need to generate cross-lane data in the created worlds; here, we utilize Carla [19], an autonomous driving simulator platform built on Unreal Engine.
3.1 Sensor Setup
To meet the needs of most NeRF and 3D-GS algorithms, the sensors used to capture data include three color cameras (i.e. ‘left-front’, ‘front’, ‘right-front’) and one 3D laser scanner. The spatial relationships between sensors and the vehicle are fixed as shown in Fig. 2. All three RGB cameras share identical intrinsic, lens parameters. Specifically, they have a resolution of $1920\times 1280$, a field of view (FOV) of $49.5^{\circ}\times 36.7^{\circ}$, a sensitivity of ISO 100, and a shutter speed of 5ms. The 3D laser scanner features 64 laser beams, a scanning range of 60 meters, and laser beam angles ranging from $-30^{\circ}$ to $30^{\circ}$ in the vertical direction.
3.2 Data Generation
Our simulation environment consists of six scenes (“Town01,” “Town02,” “Town03,” “Town04,” “Town05,” and “Town10”, all of them are under CC-BY License provided by [19]) with various weather conditions, such as sunshine and rain, closely resembling real-world settings. An overview of all the scenes and the trajectories is shown in Fig. 1. For example, the training set samples 150 times in one scene to capture the images of three cameras, LiDAR points, and the vehicle’s extrinsic information. The vehicle’s forward distance between the two sample points is 2 meters. For the evaluation process, in contrast to datasets captured in real-world scenarios, our cameras are configured with identical parameters, ensuring uniform image quality across all cameras. Therefore, we only test the novel-view-synthesis in the view of front camera, which includes four groups. These images are aligned parallel to the training set, with each group exhibiting a progressive deviation of 0m, 1m, 2m, and 4m along the y-axis in vehicle coordinate, as shown in Fig. 2. Additionally, a few novel-view-synthesis for AD methods like GaussianPro [16], MARS [57], and UC-NeRF [15] need annotated sky masks to split the scene into foreground-sky and model the color compensation separately. We employ a pre-trained SegFormer [58] to effectively infer semantic segmentation masks and extract sky masks from them.
4 Benchmark
Table 1: Results on our proposed dataset with the different offsets using front-only camera. $\uparrow$: higher is better, $\downarrow$: lower is better. The red, orange, and yellow colors respectively denote the best, the second best, and the third best results.
Method w/o Offset Offset-1m Offset-2m Offset-4m
[…]
- NeRF-based
Instant-NGP [38] 29.76 0.894 0.253 23.44 0.814 0.346 22.37 0.790 0.386 20.95 0.768 0.443
UC-NeRF [15] 35.95 0.936 0.311 30.07 0.896 0.355 25.17 0.863 0.367 22.89 0.797 0.420
MARS [57] 30.21 0.873 0.146 27.40 0.851 0.169 24.95 0.847 0.194 23.29 0.818 0.235
NeRFacto [48] 27.39 0.888 0.252 23.49 0.824 0.314 21.64 0.786 0.379 20.82 0.769 0.412
EmerNeRF [60] 31.76 0.907 0.126 28.66 0.878 0.150 26.05 0.852 0.182 24.80 0.837 0.203
- Gaussian-based
3DGS [26] 30.87 0.916 0.274 23.26 0.873 0.334 22.01 0.829 0.396 19.17 0.768 0.460
PVG [13] 37.78 0.960 0.189 26.84 0.882 0.296 24.42 0.854 0.335 23.17 0.841 0.353
GaussianPro [16] 31.62 0.919 0.263 22.61 0.856 0.338 21.26 0.819 0.383 18.75 0.772 0.445
DC-Gaussian [53] 31.29 0.919 0.264 26.82 0.884 0.298 25.24 0.871 0.319 22.90 0.844 0.360
Table 2: Results on our proposed dataset with the different offsets using left-front, front, right-front cameras. $\uparrow$: higher is better, $\downarrow$: lower is better. The red, orange, and yellow colors respectively denote the best, the second best, and the third best results.
Method w/o Offset Offset-1m Offset-2m Offset-4m
[…]
- NeRF-based
Instant-NGP [38] 29.24 0.888 0.262 23.52 0.847 0.344 22.39 0.815 0.382 21.05 0.783 0.428
UC-NeRF [15] 33.12 0.912 0.360 30.07 0.896 0.355 28.81 0.881 0.373 26.87 0.870 0.421
MARS [57] 31.37 0.887 0.151 29.28 0.874 0.157 28.54 0.869 0.165 26.49 0.847 0.193
NeRFacto [48] 29.39 0.890 0.246 22.91 0.850 0.307 22.26 0.809 0.348 21.03 0.779 0.393
EmerNeRF [60] 31.51 0.894 0.146 29.41 0.878 0.152 28.66 0.873 0.160 26.62 0.851 0.188
- Gaussian-based
3DGS [26] 29.74 0.914 0.312 22.08 0.842 0.359 21.34 0.824 0.402 19.47 0.796 0.441
PVG [13] 33.33 0.933 0.256 26.62 0.878 0.318 25.50 0.870 0.332 23.35 0.849 0.360
GaussianPro [16] 28.13 0.889 0.321 21.90 0.839 0.379 20.85 0.822 0.404 19.36 0.795 0.443
DC-Gaussian [53] 30.23 0.912 0.271 26.74 0.883 0.315 25.53 0.872 0.329 23.53 0.860 0.357
4.1 Benchmarking Environment
To comprehensively evaluate the performance and computational efficiency of the assessment methods, we conducted a series of experiments using an NVIDIA Tesla V100 16GB GPU. We benchmarked the selected methods across five different cities that include various driving scenarios. Our findings are presented through both qualitative and quantitative analyses.
Figure 3: Visualization of the rendered images using DC-Gaussian and EmerNeRF with different offsets (i.e. 0m, 2m, and 4m) in two scenes. The discriminated areas are highlighted, and the areas with better results are marked as $\square$, while worse results are marked as $\square$.
4.2 Benchmarking Methods
InstantNGP [38]: We employ the Adam optimizer and maintain similar parameter settings as the original Instant-NGP implementation: the learning rate is $1\times 10^{-4}$, the number of feature dimensions per entry is $F=8$, the number of levels is $L=10$, and the hash tables is $2^{4}$. We train the model with 30,000 steps.
Nerfacto [48]: We use the implementation in [48] without pose refinement to test our benchmark. We employ the Adam optimizer with $1\times 10^{-3}$ learning rate. We train the model with 30,000 steps.
MARS [57]: We inherit most of the parameter settings as the original MARS implementation. We employ the RAdam optimizer with $1\times 10^{-3}$ learning rate. Since our scenes are static without moving objects, we disable $\mathcal{L}_{sem}$, and the rest of the loss functions remain the same. We train the model with 50,000 steps.
UC-NeRF [15]: We inherit most parameter settings from the original UC-NeRF implementation. We employ the AdamW optimizer with $2.5\times 10^{-3}$ learning rate. The weight of sky loss is set to $2\times 10^{-3}$. We train the model with 40,000 steps.
3DGS [26]: We use the implementation of NerfStudio to evaluate our dataset. We employ the AdamW optimizer with $1\times 10^{-3}$ learning rate. For stability, we “warm up” the computation in lower resolution. Specifically, we start the optimization using 4 times smaller image resolution and then upsample twice after 500 and 1000 iterations. We train the model with 30,000 steps.
PVG [13]: We employ the Adam optimizer and maintain a similar learning rate for most parameters as the original PVG implementation. At the same time, we adjust the learning rate of the amplitude $\mathbf{A}$, opacity decaying $\mathbf{\beta}$ and opacity $\mathbf{O}$ to $3\times 10^{-5}$, $0.02$ and $0.005$ respectively. We train the model with 30,000 steps.
GaussianPro [16]: In alignment with the approach described in GaussianPro, our models are trained for 30,000 iterations across all scenes following GaussianPro’s training schedule and hyper-parameters. The interval step of the progressive propagation strategy is set to 50 where propagation is performed 3 times. The threshold $\sigma$ of the absolute relative difference is set to 0.8. We set $\beta=0.001$ and $\gamma=0.001$ for the planar loss.
DC-Gaussian [53]: To align the performance described in DC-Gaussian, we set loss coefficient 0.001 for both photometric and sky losses. Sky loss is the same as UC-NeRF.
4.3 Used Metrics
We adopt the evaluation criteria employed by the methods mentioned above, which comprise Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) as our evaluation metrics. Furthermore, we ensure transparency and clarity by describing the experimental framework employed to compare the methods.
4.4 Experimental Results
We evaluate the methods mentioned below on our dataset using novel-view-synthesis metrics. These methods are trained in two modes: (1) front-only mode (Tab.  1) utilizes images captured solely by the front camera; (2) multi-camera mode (Tab.  12) incorporates images from all three cameras. We conduct separate evaluations on trajectories with 0m, 1m, 2m, and 4m offsets. Our evaluation encompasses qualitative and quantitative experiments, ensuring a comprehensive analysis of the methods’ performance. Furthermore, it is worth noting that EmerNeRF [60] demonstrates the highest performance in the cross-lane dataset, with an average PSNR of 26.50 dB. We attribute this superior performance to its inherent self-supervised scene decomposition and positional embedding decomposition capability. We place detailed experimental results in Tab. 1 and Tab. 12. For more detailed results, please refer to the supplementary material and our webpage.
Figure 4: Visualization of the rendered images and depth maps using 3D-GS [26], GaussianPro [16], and PVG [13] with different offsets (i.e. 0m, 2m, and 4m) in two scenes. The discriminated areas are highlighted, and the areas with better results are marked as $\square$’.
5 Findings
[…]
Lately, there has been a shift in research interest within the community, transitioning from NeRF towards 3D-GS. 3D-GS has achieved state-of-the-art results on datasets such as KITTI and Waymo. However, it is worth noting that these datasets primarily focus on evaluating interpolation images without considering significant offsets (i.e. cross-lane NVS). Consequently, a crucial aspect of our dataset is to assess and compare the novel-view synthesis performance between NeRF and 3D-GS methods under challenging conditions involving substantial offsets.
As shown in Tab. 12, in the scenario with no offset, compared with EmerNeRF [60], PVG [13] achieves better performance by 0.2dB in PSNR metrics with 3-cameras setting and 1.83dB in PSNR metrics with 1-camera setting. In the scenario with 4m offset, EmerNeRF [60] and UC-NeRF [15] show a much more powerful ability to synthesize images in cross-lane novel view. compared with the SOTA Gaussian method (i.e. DC-Gaussian [53]), they achieve leading performance by 1.90dB and 3.43dB in PSNR separately. The visualization results are shown in Fig. 3.
[…]
In this comparison, we examine the self-decomposition method (PVG [13]) alongside traditional approaches (3D-GS [26] and GaussianPro [16]), the quantitative results are shown in Tab. 3. The traditional method initializes 3D-GS primitives, denoted as $\mathcal{G}=\left\{\left(\sigma_{i},\boldsymbol{\mu}_{i},\Sigma_{i},c_{i}\right)\right\}_{i=1}^{G}$, using LiDAR point clouds and extends them to areas lacking geometric features through cloning and splitting operations. However, these techniques lack a geometry prior and heavily rely on photometric loss, resulting in an issue of over-fitting. This problem becomes evident in our benchmark, particularly in the context of cross-lane novel view synthesis, as illustrated in Fig. 4.
Table 3: Results on three 3D-GS methods with the different offsets with multi-cameras setting.
Offset 3D-GS GaussianPro PVG
[…]
0m 27.23 0.892 0.278 26.97 0.890 0.283 34.08 0.948 0.219
1m 22.33 0.846 0.331 21.89 0.840 0.335 28.11 0.907 0.270
2m 21.15 0.825 0.355 20.76 0.819 0.362 26.63 0.894 0.289
4m 19.24 0.797 0.393 19.01 0.792 0.399 23.96 0.873 0.317
When not initialized by point clouds, the Gaussian points representing the background tend to overfit the training data to minimize rendering loss, often at the expense of positional accuracy. While they may appear satisfactory within the training trajectory, their performance degrades significantly when rendering images with offsets from the training trajectory. In Fig. 4, it can be observed that the street lamps are accurately modeled in the rendered images without any offset. However, when the offset is increased to 2m and 4m, these lamps become obscured by the background Gaussian points, resulting in a substantial decline in performance. Specifically, there is a decrease in performance of -7.99dB and -7.69dB for 3D-GS [26] and GaussianPro [16], respectively, as we move from 0 meters to 4 meters offset.
In contrast, the self-decomposition method, such as PVG [13], incorporates frequency information into 3D-GS primitives to create a unified representation of dynamic objects and the background. These primitives are denoted as $\mathcal{G}^{k}_{t}=\{\tilde{\mu}^{k}_{t},\Sigma^{k},\widetilde{\alpha}^{k}_{t},S^{k}\}_{k=1}^{G}$. Additionally, PVG employs a high-resolution environment cube map, denoted as $f_{sky}(d)=c_{sky}$, to effectively handle high-frequency details in the sky, where $d$ represents the ray direction. These techniques enable the network to model the sky and avoid local minima accurately. As depicted in Fig. 4, PVG produces depth maps that better represent the background, leading to more realistic images than other Gaussian methods. It achieves an impressive improvement of 4.72dB in PSNR compared to the baseline 3D-GS [26], resulting in higher fidelity rendered images.
[…]
Figure 5: Use EmerNeRF [60] as example. In the left column: we visualize the novel-view-synthesis results with different offsets and camera numbers, the discriminate areas are highlighted. In the right column, we show the performance improvement between 3-cameras and 1-camera settings using PSNR, SSIM, and LPIPS.
[…]
As shown in the right column of Fig. 6, all the selected methods perform well in the scenarios without any offset, which is identical to the previous validation efforts, and PVG even achieves leading performance up to 37.78dB in PSNR. When increasing the offset to 4 meters, its performance drops significantly while UC-NeRF [15] and EmerNeRF [60] surpass PVG by 3.16dB and 1.63dB in PSNR metric.
To this end, we visualize the rendered RGB images and depth maps with different methods, as shown in the left column of Fig. 6. UC-NeRF and EmerNeRF perform reasonable depth maps, properly distributing background buildings and sky. PVG successfully detaches the sky but performs poorly on the background buildings. GaussianPro overfits the training images and fails to estimate depth. These seriously affect the render performance in novel views with large offsets, where UC-NeRF, EmerNeRF, and PVG can render the rough structure of the scene in novel views, but GaussianPro fails. These phenomena show that the NeRF / Gaussian methods need to achieve precise geometry reconstruction to obtain better performance in cross-lane NVS.
Figure 6: In the left column: we visualize the novel-view-synthesis results of four NeRF and Gaussian methods with different offsets. In the right column, we demonstrate the quantitative comparison between these four methods with different offsets using PSNR metrics.
5.1 Discussion
The results show that current approaches exhibit a substantial gap, indicating their limited capability to meet the rigorous requirements of cross-lane or closed-loop simulation. Future research endeavors can leverage our proposed dataset and benchmark to gauge how novel methods can advance toward achieving closed-loop simulation.
Limitation.  Currently, our dataset has certain limitations as it was generated using the Unreal-based Carla simulator. We see our work represents an initial stride towards the accurate evaluation of novel driving view synthesis. As a future endeavor, we plan to curate a real-world dataset, similar to [40], encompassing cross-lane ground truth data. This expansion will enhance the authenticity and applicability of our evaluation, advancing the field of novel driving view synthesis.
6 Conclusion
In conclusion, this paper addresses the challenge that existing evaluation methods for NVS fall short of meeting the requirements of closed-loop simulations, which demand the capability to render views beyond the original trajectory. We introduce a unique driving view synthesis dataset and benchmark for autonomous driving simulations. This dataset includes testing images captured by deviating from the training trajectory, enabling the realistic evaluation of NVS approaches. Our dataset establishes a much-needed benchmark for advancing NVS techniques in closed-loop autonomous driving simulation.
Acknowledgements
The authors would like to thank Yuanyuan Gao and Jingfeng Li for their assistance in preparing the experiments and the reviewers for their insightful feedback.
Supplementary Material
Figure 7: The ground truth and rendered results, obtained from various benchmarking methods, are compared across different scenes with a 1-meter offset.
Figure 8: The ground truth and rendered results, obtained from various benchmarking methods, are compared across different scenes with a 4-meter offset.
In this supplementary, we present the evaluation results of novel view synthesis using both front-only and multi-camera settings. For the front-only camera setup, we detail the results with various offsets (e.g., 0m, 1m, 2m, 4m), as illustrated in Tables 4, 5, 6, and 7. Additionally, we provide comprehensive experimental results for the multi-camera configuration, which are shown in Table 12. The novel view synthesis results for this setting, also with different offsets (e.g., 0m, 1m, 2m, 4m), are detailed in Tables 8, 9, 10, and 11.
We also present additional results of various methods across different offsets for the Town-01 scene, as illustrated in Fig. 9. Furthermore, we showcase the visualization results for each scene using different methods. Specifically, Fig. 7 displays the results with a 1-meter offset, while Fig. 8 presents the results with a 4-meter offset.
Figure 9: The ground truth and rendered results by different benchmarking methods from a sequence under different meters of offsetting.
Table 4: Results on our proposed benchmark with the 0 meter offset with 1 camera.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 31.64 0.919 0.214 30.95 0.881 0.265 32.13 0.928 0.203
UC-NeRF [15] 39.35 0.958 0.315 35.19 0.941 0.288 34.53 0.939 0.303
MARS [57] 32.70 0.896 0.140 31.87 0.921 0.100 31.76 0.894 0.125
NeRFacto [48] 28.74 0.905 0.253 28.40 0.901 0.224 31.25 0.924 0.190
EmerNeRF [60] 33.43 0.932 0.111 31.72 0.914 0.106 32.33 0.922 0.102
- Gaussian-based
3D-GS [26] 32.12 0.937 0.326 29.09 0.908 0.295 29.62 0.916 0.248
PVG [13] 39.16 0.971 0.212 38.81 0.973 0.12 38.08 0.968 0.171
GaussianPro [16] 35.51 0.953 0.266 30.56 0.915 0.251 29.39 0.910 0.242
DC-Gaussian [53] 35.51 0.953 0.266 30.56 0.915 0.251 29.39 0.91 0.242
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 29.58 0.921 0.221 27.37 0.892 0.241 26.86 0.820 0.372
UC-NeRF [15] 39.03 0.957 0.300 33.60 0.923 0.321 34.00 0.900 0.340
MARS [57] 36.68 0.957 0.084 26.31 0.857 0.188 30.44 0.875 0.151
NeRFacto [48] 26.09 0.899 0.265 23.64 0.871 0.261 26.23 0.826 0.316
EmerNeRF [60] 36.42 0.944 0.094 26.63 0.873 0.175 30.03 0.854 0.167
- Gaussian-based
3D-GS [26] 33.60 0.946 0.280 25.95 0.872 0.380 28.45 0.841 0.390
PVG [13] 39.58 0.971 0.194 34.91 0.954 0.194 36.16 0.921 0.244
GaussianPro [16] 33.69 0.943 0.273 30.54 0.916 0.249 30.03 0.879 0.301
DC-Gaussian [53] 33.69 0.943 0.273 26.17 0.916 0.249 30.03 0.879 0.301
Table 5: Results on our proposed benchmark with the 1-meter offset with 1 camera.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 24.19 0.868 0.350 20.25 0.766 0.390 23.92 0.841 0.288
UC-NeRF [15] 29.60 0.925 0.355 28.68 0.894 0.338 30.11 0.911 0.336
MARS [57] 28.27 0.886 0.145 27.95 0.883 0.138 29.04 0.899 0.140
NeRFacto [48] 24.48 0.884 0.300 24.40 0.842 0.279 23.55 0.835 0.278
EmerNeRF [60] 28.81 0.913 0.123 27.82 0.876 0.143 29.39 0.901 0.126
- Gaussian-based
3D-GS [26] 24.07 0.909 0.321 19.98 0.799 0.372 21.22 0.811 0.334
PVG [13] 25.62 0.917 0.282 27.25 0.912 0.267 28.15 0.907 0.253
GaussianPro [16] 23.97 0.902 0.329 20.71 0.811 0.360 21.92 0.840 0.322
DC-Gaussian [53] 27.34 0.923 0.272 24.87 0.873 0.303 26.35 0.891 0.275
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 26.21 0.869 0.277 22.18 0.800 0.338 23.88 0.739 0.434
UC-NeRF [15] 34.10 0.938 0.338 29.42 0.890 0.353 28.51 0.820 0.410
MARS [57] 34.28 0.937 0.097 24.42 0.831 0.210 27.25 0.812 0.199
NeRFacto [48] 24.20 0.858 0.311 21.00 0.794 0.336 23.30 0.732 0.380
EmerNeRF [60] 34.15 0.930 0.102 24.83 0.851 0.194 26.96 0.797 0.211
- Gaussian-based
3D-GS [26] 25.62 0.905 0.343 21.52 0.840 0.346 21.75 0.843 0.342
PVG [13] 29.19 0.926 0.286 25.33 0.864 0.293 25.51 0.764 0.397
GaussianPro [16] 25.56 0.907 0.341 21.52 0.840 0.346 21.92 0.841 0.335
DC-Gaussian [53] 30.67 0.936 0.266 26.17 0.883 0.297 25.51 0.796 0.376
Table 6: Results on our proposed benchmark with the 2-meter offset with 1 camera.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 24.21 0.861 0.377 20.14 0.767 0.405 21.61 0.802 0.344
UC-NeRF [15] 28.45 0.917 0.373 27.53 0.895 0.347 29.04 0.908 0.343
MARS [57] 26.58 0.875 0.165 27.29 0.886 0.132 27.37 0.886 0.151
NeRFacto [48] 23.73 0.869 0.334 22.95 0.818 0.312 21.29 0.795 0.333
EmerNeRF [60] 27.12 0.902 0.143 27.16 0.879 0.137 27.72 0.888 0.137
- Gaussian-based
3D-GS [26] 22.00 0.881 0.369 19.28 0.796 0.394 21.91 0.840 0.322
PVG [13] 24.58 0.910 0.306 26.13 0.885 0.293 26.58 0.894 0.272
GaussianPro [16] 21.89 0.879 0.368 19.28 0.796 0.394 21.91 0.84 0.322
DC-Gaussian [53] 25.89 0.912 0.294 23.28 0.866 0.314 24.61 0.876 0.299
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 24.71 0.841 0.327 21.03 0.776 0.387 22.5 0.694 0.474
UC-NeRF [15] 31.31 0.925 0.353 28.46 0.882 0.363 26.95 0.907 0.419
MARS [57] 31.57 0.924 0.120 24.19 0.819 0.222 26.12 0.795 0.214
NeRFacto [48] 24.54 0.873 0.301 20.25 0.772 0.381 21.97 0.685 0.419
EmerNeRF [60] 31.44 0.917 0.125 24.60 0.839 0.206 25.83 0.780 0.226
- Gaussian-based
3D-GS [26] 23.35 0.878 0.384 20.52 0.817 0.381 20.53 0.806 0.458
PVG [13] 28.27 0.914 0.312 24.41 0.853 0.311 24.13 0.744 0.411
GaussianPro [16] 23.54 0.881 0.381 20.52 0.817 0.381 20.45 0.706 0.457
DC-Gaussian [53] 28.27 0.920 0.294 25.21 0.879 0.321 24.19 0.773 0.392
Table 7: Results on our proposed benchmark with the 4-meter offset with 1 camera.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 23.20 0.854 0.401 19.44 0.754 0.477 19.49 0.764 0.413
UC-NeRF [15] 26.12 0.903 0.410 24.34 0.872 0.400 27.23 0.888 0.373
MARS [57] 23.45 0.854 0.202 24.41 0.848 0.177 25.00 0.843 0.196
NeRFacto [48] 21.70 0.847 0.366 20.14 0.767 0.405 19.15 0.756 0.402
EmerNeRF [60] 23.89 0.876 0.184 24.31 0.843 0.181 25.39 0.862 0.180
- Gaussian-based
3D-GS [26] 19.57 0.842 0.424 17.42 0.755 0.449 18.01 0.773 0.414
PVG [13] 22.59 0.893 0.336 22.74 0.862 0.327 23.65 0.866 0.310
GaussianPro [16] 19.50 0.840 0.420 17.39 0.753 0.447 17.96 0.767 0.411
DC-Gaussian [53] 23.56 0.897 0.329 22.35 0.835 0.362 22.06 0.854 0.338
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 22.59 0.817 0.409 19.95 0.764 0.422 21.02 0.652 0.533
UC-NeRF [15] 27.96 0.890 0.409 26.27 0.852 0.397 26.07 0.755 0.473
MARS [57] 27.39 0.888 0.175 23.37 0.813 0.239 24.36 0.749 0.258
NeRFacto [48] 20.90 0.815 0.419 19.49 0.760 0.414 20.39 0.637 0.489
EmerNeRF [60] 27.35 0.886 0.177 23.66 0.814 0.227 24.17 0.739 0.266
- Gaussian-based
3D-GS [26] 20.39 0.835 0.451 19.01 0.795 0.426 18.42 0.648 0.515
PVG [13] 25.11 0.891 0.359 22.90 0.831 0.338 22.00 0.700 0.447
GaussianPro [16] 20.39 0.835 0.451 18.91 0.783 0.43 18.33 0.647 0.511
DC-Gaussian [53] 24.26 0.893 0.345 23.18 0.850 0.356 21.99 0.737 0.427
Table 8: Results on our proposed benchmark with the 0-meter offset with 3 cameras.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 29.56 0.906 0.250 30.95 0.907 0.236 31.13 0.911 0.217
UC-NeRF [15] 35.43 0.936 0.375 31.40 0.904 0.365 31.95 0.921 0.344
MARS [57] 32.52 0.890 0.154 30.99 0.903 0.133 31.99 0.889 0.136
NeRFacto [48] 30.75 0.902 0.198 29.64 0.894 0.238 30.83 0.907 0.214
EmerNeRF [60] 33.25 0.926 0.125 30.84 0.896 0.139 32.56 0.917 0.113
- Gaussian-based
3D-GS [26] 31.35 0.935 0.310 27.09 0.883 0.315 27.23 0.892 0.278
PVG [13] 34.59 0.955 0.258 33.77 0.948 0.223 34.08 0.948 0.219
GaussianPro [16] 31.07 0.934 0.211 27.23 0.883 0.215 26.97 0.888 0.283
DC-Gaussian [53] 33.95 0.948 0.258 29.19 0.910 0.262 28.58 0.914 0.251
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 29.73 0.899 0.222 26.69 0.871 0.278 27.39 0.836 0.366
UC-NeRF [15] 36.29 0.942 0.332 32.51 0.909 0.340 31.15 0.859 0.403
MARS [57] 35.72 0.942 0.111 28.28 0.856 0.170 28.73 0.843 0.203
NeRFacto [48] 29.65 0.887 0.268 27.46 0.898 0.277 28.00 0.853 0.281
EmerNeRF [60] 35.46 0.929 0.121 28.60 0.872 0.157 28.32 0.822 0.219
- Gaussian-based
3D-GS [26] 30.62 0.924 0.320 31.25 0.930 0.321 30.91 0.922 0.326
PVG [13] 35.41 0.956 0.243 30.79 0.916 0.263 31.36 0.874 0.329
GaussianPro [16] 30.45 0.923 0.322 25.69 0.871 0.328 27.39 0.836 0.366
DC-Gaussian [53] 32.74 0.947 0.251 27.91 0.898 0.275 29.01 0.857 0.327
Table 9: Results on our proposed benchmark with the 1-meter offset with 3 cameras.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 22.51 0.881 0.361 20.41 0.794 0.358 23.62 0.892 0.263
UC-NeRF [15] 29.60 0.925 0.355 28.68 0.894 0.338 30.11 0.911 0.336
MARS [57] 28.58 0.888 0.149 28.81 0.884 0.144 30.10 0.907 0.129
NeRFacto [48] 23.60 0.891 0.311 22.70 0.862 0.272 21.98 0.826 0.286
EmerNeRF [60] 29.12 0.915 0.127 28.68 0.877 0.149 30.448 0.909 0.115
- Gaussian-based
3D-GS [26] 23.90 0.905 0.340 20.67 0.818 0.382 22.33 0.846 0.331
PVG [13] 25.48 0.92 0.290 25.82 0.884 0.295 28.11 0.907 0.270
GaussianPro [16] 23.37 0.902 0.349 20.81 0.817 0.353 22.19 0.0.740 0.424
DC-Gaussian [53] 28.01 0.928 0.288 25.08 0.869 0.315 25.92 0.891 0.289
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 26.60 0.897 0.287 23.15 0.854 0.321 24.81 0.763 0.472
UC-NeRF [15] 34.10 0.938 0.338 29.42 0.890 0.353 28.51 0.820 0.410
MARS [57] 33.96 0.930 0.110 26.42 0.832 0.190 27.84 0.807 0.220
NeRFacto [48] 24.08 0.869 0.300 20.50 0.793 0.351 24.58 0.861 0.319
EmerNeRF [60] 33.83 0.923 0.115 26.83 0.852 0.174 27.55 0.792 0.232
- Gaussian-based
3D-GS [26] 23.19 0.896 0.377 19.66 0.827 0.385 22.72 0.757 0.336
PVG [13] 29.08 0.926 0.305 25.54 0.864 0.319 25.69 0.769 0.428
GaussianPro [16] 23.45 0.896 0.379 19.50 0.819 0.389 22.69 0.757 0.443
DC-Gaussian [53] 30.27 0.935 0.289 25.85 0.878 0.306 25.33 0.794 0.402
Table 10: Results on our proposed benchmark with the 2-meter offset with 3 cameras.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 22.51 0.881 0.361 20.41 0.794 0.358 23.62 0.892 0.263
UC-NeRF [15] 29.60 0.925 0.355 28.68 0.894 0.338 30.11 0.911 0.336
MARS [57] 28.09 0.882 0.161 28.69 0.892 0.132 29.49 0.904 0.131
NeRFacto [48] 23.57 0.842 0.323 19.97 0.769 0.393 23.86 0.728 0.408
EmerNeRF [60] 28.63 0.909 0.139 28.56 0.885 0.137 29.84 0.906 0.117
- Gaussian-based
3D-GS [26] 22.43 0.891 0.369 22.21 0.812 0.397 21.15 0.825 0.355
PVG [13] 24.31 0.912 0.313 24.74 0.885 0.298 26.63 0.894 0.289
GaussianPro [16] 22.39 0.891 0.368 20.06 0.812 0.396 20.76 0.819 0.362
DC-Gaussian [53] 26.98 0.921 0.303 24.02 0.871 0.315 24.51 0.878 0.306
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 26.60 0.897 0.287 23.15 0.854 0.321 24.81 0.763 0.472
UC-NeRF [15] 34.10 0.938 0.338 29.42 0.890 0.353 28.51 0.820 0.410
MARS [57] 31.95 0.919 0.131 25.77 0.821 0.205 27.22 0.798 0.228
NeRFacto [48] 23.42 0.878 0.337 22.52 0.853 0.282 20.22 0.785 0.343
EmerNeRF [60] 31.82 0.912 0.136 26.18 0.841 0.189 26.93 0.783 0.240
- Gaussian-based
3D-GS [26] 21.80 0.876 0.410 19.11 0.810 0.412 21.33 0.729 0.468
PVG [13] 28.10 0.917 0.325 24.65 0.855 0.333 24.58 0.756 0.435
GaussianPro [16] 21.69 0.876 0.408 18.70 0.804 0.426 21.48 0.732 0.465
DC-Gaussian [53] 28.25 0.919 0.313 25.12 0.862 0.324 24.32 0.778 0.411
Table 11: Results on our proposed benchmark with the 4-meter offset with 3 cameras.
Scene001 Scene002 Scene003
[…]
- NeRF-based
Instant-NGP [38] 22.04 0.866 0.414 19.95 0.781 0.405 18.33 0.745 0.412
UC-NeRF [15] 26.75 0.904 0.422 26.09 0.867 0.413 27.66 0.887 0.393
MARS [57] 25.53 0.871 0.185 26.97 0.868 0.158 28.04 0.870 0.155
NeRFacto [48] 22.41 0.862 0.367 21.51 0.817 0.326 18.33 0.745 0.412
EmerNeRF [60] 25.97 0.893 0.167 26.87 0.863 0.162 28.43 0.889 0.139
- Gaussian-based
3D-GS [26] 20.73 0.872 0.402 18.86 0.786 0.436 19.24 0.797 0.393
PVG [13] 22.40 0.898 0.338 22.56 0.860 0.331 23.96 0.873 0.317
GaussianPro [16] 22.04 0.866 0.414 19.95 0.781 0.405 18.33 0.745 0.412
DC-Gaussian [53] 25.00 0.908 0.329 22.07 0.851 0.347 22.61 0.867 0.333
Scene004 Scene005 Scene006
[…]
- NeRF-based
Instant-NGP [38] 23.23 0.835 0.380 19.70 0.771 0.434 23.03 0.700 0.522
UC-NeRF [15] 27.83 0.886 0.416 26.79 0.852 0.399 26.11 0.759 0.482
MARS [57] 28.38 0.888 0.177 23.90 0.818 0.227 26.10 0.768 0.256
NeRFacto [48] 22.38 0.815 0.382 19.04 0.756 0.427 22.53 0.681 0.446
EmerNeRF [60] 28.34 0.886 0.179 24.19 0.819 0.215 25.91 0.758 0.264
- Gaussian-based
3D-GS [26] 20.31 0.845 0.460 17.90 0.785 0.453 19.87 0.693 0.499
PVG [13] 25.11 0.897 0.365 23.14 0.840 0.352 22.95 0.724 0.458
GaussianPro [16] 23.23 0.835 0.380 19.70 0.771 0.434 23.03 0.700 0.522
DC-Gaussian [53] 23.35 0.848 0.348 25.12 0.896 0.355 23.02 0.788 0.431
Table 12: Results on our proposed dataset with the different offsets using left-front, front, right-front cameras.
Method w/o Offset Offset-1m Offset-2m Offset-4m
[…]
- NeRF-based
Instant-NGP [38] 29.24 0.888 0.262 23.52 0.847 0.344 22.39 0.815 0.382 21.05 0.783 0.428
UC-NeRF [15] 33.12 0.912 0.360 30.07 0.896 0.355 28.81 0.881 0.373 26.87 0.870 0.421
MARS [57] 31.37 0.887 0.151 29.28 0.874 0.157 28.54 0.869 0.165 26.49 0.847 0.193
NeRFacto [48] 29.39 0.890 0.246 22.91 0.850 0.307 22.26 0.809 0.348 21.03 0.779 0.393
EmerNeRF [60] 31.51 0.894 0.146 29.41 0.878 0.152 28.66 0.873 0.160 26.62 0.851 0.188
- Gaussian-based
3DGS [26] 29.74 0.914 0.312 22.08 0.842 0.359 21.34 0.824 0.402 19.47 0.796 0.441
PVG [13] 33.33 0.933 0.256 26.62 0.878 0.318 25.50 0.870 0.332 23.35 0.849 0.360
GaussianPro [16] 28.13 0.889 0.321 21.90 0.839 0.379 20.85 0.822 0.404 19.36 0.795 0.443
DC-Gaussian [53] 30.23 0.912 0.271 26.74 0.883 0.315 25.53 0.8715 0.329 23.53 0.860 0.357